12) Random variables X & Y have joint pmf given in the table.

Y = 1 Y = 2 Y= 3

X = 1 0.3 0.1 0

X = 2 0.1 0.3 0.2

In problem (12), determine E(3Y + 1.2 | X < Y ) a) 9.0 b) 9.2 c) 9.5 d) 9.8 e) 10.0
Answer #1

here P(X<Y)=P(X=1,Y=2)+P(X=1,Y=3)+P(X=2,Y=3)=0.1+0+0.2=0.3

hence E( 3Y + 1.2 | X < Y ) =(3*2+1.2)*(0.1/0.3)+(3*3+1.2)*(0.2/0.3)=9.2

option B is correct